# Stationary Policies are Optimal in Risk-averse Total-reward MDPs with EVaR

### v1

@@ Abstract
Optimizing risk-averse objectives in discounted MDPs is challenging because most models do not admit direct dynamic programming equations and require complex history-dependent policies. In this paper, we show that the risk-averse total reward criterion, under the Entropic Risk Measure (ERM) and Entropic Value at Risk (EVaR) risk measures, can be optimized by a stationary policy, making it simple to analyze, interpret, and deploy. We propose exponential value iteration, policy iteration, and linear programming to compute optimal policies. In comparison with prior work, our results only require the relatively mild condition of transient MDPs and allow for both positive and negative rewards. Our results indicate that the total reward criterion may be preferable to the discounted criterion in a broad range of risk-averse reinforcement learning domains.
@@ 1 Introduction
Risk-averse Markov decision processes (MDP) (Puterman 2005) that use monetary risk measures as their objective have been gaining in popularity in recent years (Kastner, Erdogdu, and Farahmand 2023; Marthe, Garivier, and Vernade 2023; Lam et al. 2022; Li, Zhong, and Brandeau 2022; Bäuerle and Glauner 2022; Hau, Petrik, and Ghavamzadeh 2023; Hau et al. 2023; Su, Petrik, and Grand-Clément 2024). Risk-averse objectives, such as Value-at-Risk (VaR) or Conditional-Value-at-Risk (CVaR) penalize the variability of returns (Follmer and Schied 2016). As a result, one prefers policies with stronger guarantees on the probability of catastrophic losses. These objectives are important in critical domains, such as healthcare or finance, where one must avoid catastrophic failures.
@@
In a departure from much of recent literature, we target the total reward criterion (TRC) (Kallenberg 2021; Puterman 2005) instead of the common discounted criterion. TRC, which generalizes the stochastic shortest path problem, assumes an infinite horizon but does not discount future rewards. To control for infinite returns, the TRC assumes a positive probability that the process terminates.
@@
Two essential reasons motivate our departure from discounted objectives in risk-averse MDPs. First, computing discounted risk-averse policies is difficult. Reinforcement learning algorithms typically use discounted objectives because they admit optimal stationary policies and value functions that can be computed using dynamic programs. In contrast, most risk measures, such as VaR or CVaR, require that optimal policies are history-dependent (Bäuerle and Ott 2011; Hau et al. 2023). The simplest known risk-averse optimal policies are Markov (time-dependent) policies for objectives defined using Entropic Value-at-Risk (EVaR) and Entropic Risk Measure (ERM) risk measures (Hau, Petrik, and Ghavamzadeh 2023).
@@
Second, TRC captures the concept of stochastic termination, which is common in reinforcement learning (Sutton and Barto 2018). In risk-neutral objectives, discounting can serve well to model the probability of termination because it guarantees the same optimal policies (Puterman 2005). However, as we show in this work, no such correspondence exists with risk-averse objectives, and the difference between them may be arbitrarily significant. Modeling stochastic termination using a discount factor in risk-averse objectives is inappropriate and leads to dramatically different optimal policies.
@@
As our main contribution, we show that the risk-averse TRC with ERM and EVaR risk measures admit optimal stationary policies and optimal value functions in transient MDPs. Transient MDPs must terminate with a positive probability in a sufficiently large number of steps and are a common model used to guarantee that TRC returns are bounded. We also show that the optimal value function satisfies dynamic programming equations and show how to compute it using value iteration, policy iteration, or linear programming algorithms. These algorithms are simple and closely resemble the algorithms for solving MDPs. In summary, our results show that the risk-averse TRC criterion is easier to optimize than risk-averse discounted objectives.
@@
Our results also indicate that EVaR is a particularly interesting risk measure in reinforcement learning. ERM and the closely related exponential utility functions have been popular in sequential decision-making problems because they admit dynamic programming decompositions (Patek and Bertsekas 1999; de Freitas, Freire, and Delgado 2020; Smith and Chapman 2023; Denardo and Rothblum 1979; Hau, Petrik, and Ghavamzadeh 2023; Hau et al. 2023). Unfortunately, ERM is difficult to interpret, scale-dependent, and incomparable with popular risk measures like VaR and CVaR. Because EVaR reduces to an optimization over ERM, it preserves most of the computational advantages of ERM. Because EVaR closely approximates CVaR and VaR at the same risk level, its value is also much easier to interpret. Finally, EVaR is also a coherent risk measure, unlike ERM (Ahmadi-Javid 2012; Ahmadi-Javid and Pichler 2017).
@@
Table 1: Structure of optimal policies in risk-averse MDPs: “S”, “M” and “H” refer to Stationary, Markov and History-dependent policies respectively. Our contributions are highlighted in bold.
A recent line of work has studied the TRC with nested risk measures, also known as Markov risk measures (Ahmadi et al. 2021a). Nested risk measures apply the risk in each level of the dynamic program independently and preserve most of the favorable computational properties of risk-neutral MDPs. For example, NCVaR refers to a nested CVaR risk measure. Unfortunately, nested risk measures are difficult to interpret, and their value depends on the sequence in which the rewards are obtained in a complex and unpredictable way (Kupper and Schachermayer 2006). In addition, we show that with negative rewards, the total reward may be unbounded even when an MDP is transient.
Table 1 puts our contribution in the context of other work on risk-averse MDP objectives. While we are not aware of prior work on the TRC objective with ERM or EVaR risk-aversion, the ERM risk measure is closely related to exponential utility functions. All prior work on TRC with exponential utility functions impose some constraints on the sign of the instantaneous rewards, such as positive rewards (Blackwell 1967) or negative rewards (Bertsekas and Tsitsiklis 1991; Freire and Delgado 2016; Carpin, Chow, and Pavone 2016; de Freitas, Freire, and Delgado 2020; Fei et al. 2021; Fei, Yang, and Wang 2021; Ahmadi et al. 2021a; Cohen et al. 2021; Meggendorfer 2022). This assumptuin limits the modeling power of prior work since assuming positive (or negative) rewards means that all states are more desirable (or detrimental) than the terminal state. Several technical hurdles arise when no sign constraints exist: zero-reward cycles may exist, optimality equations may have multiple solutions, and value and policy iterations may not converge. Our novel analysis circumvents these issues by employing a novel squeeze argument that uses the monotonicity of an exponential Bellman operator and transience of MDPs.
Notation. We use a tilde to mark random variables, e.g. $\tilde{x}$. Bold lower-case letters represent vectors, and upper-case bold letters represent matrices. Sets are either calligraphic or upper-case Greek letters. The symbol $\mathbb{X}$ represents the space of real-valued random variables. When a function is defined over an index set, such as $z\colon\{1,2,\dots,N\}\to\mathbb{R}$, we also treat it interchangeably as a vector $\bm{z}\in\mathbb{R}^{n}$ such that $z_{i}=z(i),\forall i=1,\dots,n$.
@@ 2 Background on risk-averse MDPs
We focus on solving Markov decision processes (MDPs) (Puterman 2005; Su and Petrik 2023), modeled by a tuple $(\bar{\mathcal{S}},\mathcal{A},\bar{p},\bar{r},\bm{\bar{\mu}})$, where $\bar{\mathcal{S}}=\{1,2,\dots,S,S+1\}$ is the finite set of states and $\mathcal{A}=\{1,2,\ldots,A\}$ is the finite set of actions. The transition function $\bar{p}\colon\bar{\mathcal{S}}\times\mathcal{A}\to\Delta_{\bar{\mathcal{S}}}$ represents the probability $\bar{p}(s,a,s^{\prime})$ of transitioning to $s^{\prime}\in\bar{\mathcal{S}}$ after taking $a\in\mathcal{A}$ in $s\in\bar{\mathcal{S}}$ and $\bm{\bar{p}}_{sa}\in\Delta_{\bar{\mathcal{S}}}$ is such that $(\bm{\bar{p}}_{sa})_{s^{\prime}}=\bar{p}(s,a,s^{\prime})$. The function $\bar{r}\colon\bar{\mathcal{S}}\times\mathcal{A}\times\bar{\mathcal{S}}\to\mathbb{R}$ represents the reward $\bar{r}(s,a,s^{\prime})\in\mathbb{R}$ associated with transitioning from $s\in\bar{\mathcal{S}}$ and $a\in\mathcal{A}$ to $s^{\prime}\in\bar{\mathcal{S}}$. The vector $\bm{\bar{\mu}}\in\Delta_{\bar{\mathcal{S}}}$ is the initial state distribution.
@@
The solution to an MDP is a policy. Given a horizon $t\in\mathbb{N}$, a history-dependent policy in the set $\Pi_{\mathrm{HR}}^{t}$ maps the history of states and actions to a distribution over actions. A Markov policy $\pi\in\Pi_{\mathrm{MR}}^{t}$ is a sequence of decision rules $\pi=(\bm{d}_{0},\bm{d}_{1},\dots,\bm{d}_{t-1})$ with $\bm{d}_{k}\colon\mathcal{S}\to\Delta_{\mathcal{A}}$ the decision rule for taking actions at time $k$. The set of all randomized decision rules is $\mathcal{D}=(\Delta_{\mathcal{A}})^{\mathcal{S}}$. Stationary policies $\Pi_{\mathrm{SR}}$ are of Markov policies with $\pi:=(\bm{d})_{\infty}:=(\bm{d},\bm{d},\dots)$ with the identical decision rule in every timestep. We treat decision rules and stationary policies interchangeably. The sets of deterministic Markov and stationary policies are denoted by $\Pi_{\mathrm{MD}}^{t}$ and $\Pi_{\mathrm{SD}}$. Finally, we omit the superscript $t$ to indicate infinite horizon definitions of policies.
@@
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\mathbb{E}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right],$ (1)
@@
Unlike the discounted criterion, the TRC may be unbounded, optimal policies may not exist, or may be non-stationary (Bertsekas and Yu 2013; James and Collins 2006). We assume that all policies have a positive probability of eventually transitioning to the sink state.
@@
2.2 underlies most of our results. Transient MDPs are important because their optimal policies exist and can be chosen to be stationary deterministic (Kallenberg 2021, theorem 4.12). Transient MDPs are also common in stochastic games (Filar and Vrieze 2012) and generalize the stochastic shortest path problem (Bertsekas and Yu 2013).
An important tool in their analysis is the spectral radius $\rho\colon\mathbb{R}^{n\times n}\to\mathbb{R}$ which is defined for each $\bm{A}\in\mathbb{R}^{n\times n}$ as the maximum absolute eigenvalue: $\rho(\bm{A}):=\max_{i=1,\dots,n}|\lambda_{i}|$ where $\lambda_{i}$ is the $i$-th eigenvalue (Horn and Johnson 2013).
@@
An MDP is transient if and only if $\rho(\bm{P}^{\pi})<1$ for all $\pi\in\Pi_{\mathrm{SD}}$.
Now, let us understand the basic setting differences between a discounted MDP and a transient MDP, which are useful in demonstrating the behavior of risk-averse objectives. Consider the MDPs in Figure 1. There are one non-sink state $s$ and one action $a$. A triple tuple represents an action, transition probability, and a reward separately. Note that every discounted MDP can be converted to a transient MDP by (20) in Appendix B. For the discounted MDP, the discount factor is $\gamma$. For the transient MDP, $e$ is the sink state, and there is the probability $1-\epsilon$ of transiting from state $s$ to state $e$. Once the agent reaches the state $e$, it stays in $e$. For the risk-neutral objective, if $\gamma$ equals $\epsilon$, their value functions have identical values. Please see Proposition B.1 in Appendix B for details. However, for risk-aversion objectives, such as ERM, the value functions in a discounted MDP can diverge from those in a transient MDP.
@@
Monetary risk measures aim to generalize the expectation operator to account for the spread of the random variable. Entropic risk measure (ERM) is a popular risk measure, defined for any risk level $\beta>0$ and $\tilde{x}\in\mathbb{X}$ as (Follmer and Schied 2016)
@@
and extended to $\beta\in[0,\infty]$ as $\operatorname{ERM}_{0}[\tilde{x}]=\lim_{\beta\to 0^{+}}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\mathbb{E}[\tilde{x}]$ and $\operatorname{ERM}_{\infty}[\tilde{x}]=\lim_{\beta\to\infty}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$. ERM plays a special role in sequential decision-making because it is the only law-invariant risk measure that satisfies the tower property (Kupper and Schachermayer 2006; Marthe, Garivier, and Vernade 2023), which is essential in constructing dynamic programs (Hau, Petrik, and Ghavamzadeh 2023). Unfortunately, two major limitations of ERM hinder its practical applications. First, it is not positively homogenous and, therefore, the risk value depends on the scale of the rewards, and ERM is not coherent (Follmer and Schied 2016; Hau, Petrik, and Ghavamzadeh 2023; Ahmadi-Javid 2012). Second, the risk parameter $\beta$ is difficult to interpret and does not relate well to other common risk measures, like VaR or CVaR.
@@
and is extended to $\operatorname{EVaR}_{0}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$ and $\operatorname{EVaR}_{1}\left[\tilde{x}\right]=\mathbb{E}\left[\tilde{x}\right]$ (Ahmadi-Javid 2012). It is important to note that the supremum in (4) may not be attained even when $\tilde{x}$ is a finite discrete random variable (Ahmadi-Javid and Pichler 2017).
EVaR addresses the limitations of ERM while preserving its benefits. EVaR is coherent and positively homogenous. EVaR is also a good approximation to interpretable quantile-based risk measures, like VaR and CVaR (Ahmadi-Javid 2012; Hau, Petrik, and Ghavamzadeh 2023).
@@
Risk-averse MDPs with VaR and CVaR combined with discounted objectives have received abundant attention (Hau et al. 2023; Bäuerle and Ott 2011; Bäuerle and Glauner 2022; Pflug and Pichler 2016; Li, Zhong, and Brandeau 2022), showing that these objectives require history-dependent optimal policies. Nested risk measures with the TRC criterion may admit stationary policies that can be computed using dynamic programming (Ahmadi et al. 2021a; Meggendorfer 2022; de Freitas, Freire, and Delgado 2020; Gavriel, Hanasusanto, and Kuhn 2012). However, the TRC with nested CVaR can be unbounded, as shown in Proposition C.1, which is in contradiction to some existing results from the literature, such as theorem 1 in (Ahmadi et al. 2021a). Recent works have shown that optimal policies for discounted objective with EVaR risk measure can be chosen to be Markov and can be computed via dynamic programming (Hau, Petrik, and Ghavamzadeh 2023), building upon similar results established for ERM (Chung and Sobel 1987). However, in ERM-TRC, the value functions may also be unbounded, as shown in Proposition D.1.

### v2

# Risk-averse Total-reward MDPs with ERM and EVaR

@@ Abstract
Optimizing risk-averse objectives in discounted MDPs is challenging because most models do not admit direct dynamic programming equations and require complex history-dependent policies. In this paper, we show that the risk-averse total reward criterion, under the Entropic Risk Measure (ERM) and Entropic Value at Risk (EVaR) risk measures, can be optimized by a stationary policy, making it simple to analyze, interpret, and deploy. We propose exponential value iteration, policy iteration, and linear programming to compute optimal policies. Compared with prior work, our results only require the relatively mild condition of transient MDPs and allow for both positive and negative rewards. Our results indicate that the total reward criterion may be preferable to the discounted criterion in a broad range of risk-averse reinforcement learning domains.
@@ 1 Introduction
Risk-averse Markov decision processes (MDP) Puterman (2005) that use monetary risk measures as their objective have been gaining in popularity in recent years Kastner et al. (2023); Marthe et al. (2023); Lam et al. (2022); Li et al. (2022); Bäuerle and Glauner (2022); Hau et al. (2023b, a); Su et al. (2024c, b). Risk-averse objectives, such as Value at Risk (VaR), Conditional Value at Risk (CVaR), Entropic Risk Measure (ERM), or Entropic Value at Risk (EVaR), penalize the variability of returns Follmer and Schied (2016). As a result, these risk measures yield policies with stronger guarantees on the probability of catastrophic losses, which is important in domains like healthcare or finance.
@@
In this paper, we target the total reward criterion (TRC) Kallenberg (2021); Puterman (2005) instead of the common discounted criterion. TRC also assumes an infinite horizon but does not discount future rewards. To control for infinite returns, we assume that the MDP is transient, i.e. that there is a positive probability that the process terminates after a finite number of steps, an assumption commonly used in the TRC literature Filar and Vrieze (2012). We consider the TRC with both positive and negative rewards. When the rewards are non-positive, the TRC is equivalent to the stochastic shortest path problem, and when they are non-negative, it is equivalent to the stochastic longest path Dann et al. (2023).
@@
Two reasons motivate our departure from discounted objectives in risk-averse MDPs. First, considering risk affects discounted objectives significantly. It is common to use discounted objectives because they admit optimal stationary policies and value functions that can be computed using dynamic programs. However, most risk-averse discount objectives, such as VaR, CVaR, or EVaR, require that optimal policies are history-dependent Bäuerle and Ott (2011); Hau et al. (2023a, b) and do not admit standard dynamic programming optimality equations.
@@
Second, TRC captures the concept of stochastic termination, which is common in reinforcement learning Sutton and Barto (2018). In risk-neutral objectives, discounting can serve well to model the probability of termination because it guarantees the same optimal policies Puterman (2005); Su and Petrik (2023). However, as we show in this work, no such correspondence exists with risk-averse objectives, and the difference between them may be arbitrarily significant. Modeling stochastic termination using a discount factor in risk-averse objectives is inappropriate and leads to dramatically different optimal policies.
@@
As our main contribution, we show that the risk-averse TRC with ERM and EVaR risk measures admit optimal stationary policies and optimal value functions in transient MDPs. We also show that the optimal value function satisfies dynamic programming equations and can be computed with exponential value iteration, policy iteration, or linear programming algorithms. These algorithms are simple and closely resemble the algorithms for solving MDPs.
@@
Our results indicate that EVaR is a particularly interesting risk measure in reinforcement learning. ERM and the closely related exponential utility functions have been popular in sequential decision-making problems because they admit dynamic programming decompositions Patek and Bertsekas (1999); de Freitas et al. (2020); Smith and Chapman (2023); Denardo and Rothblum (1979); Hau et al. (2023b, a). Unfortunately, ERM is difficult to interpret; it is scale-dependent; and it is incomparable with popular risk measures like VaR and CVaR. Because EVaR reduces to an optimization over ERM, it preserves most of the computational advantages of ERM, and since EVaR closely approximates CVaR and VaR at the same risk level, its value is also much easier to interpret. Finally, EVaR is also a coherent risk measure, unlike ERM Ahmadi-Javid (2012); Ahmadi-Javid and Pichler (2017).
@@
Table 1: Structure of optimal policies in risk-averse MDPs: “S”, “M” and “H” refer to Stationary, Markov and History-dependent policies respectively.
Table 1 puts our contribution in the context of other work on risk-averse MDP objectives. Optimal policies for VaR and CVaR are known to be history-dependent in the discounted objective Bäuerle and Ott (2011); Hau et al. (2023a) and must be history-dependent in TRC because TRC generalizes the finite-horizon objective. The TRC with Nested risk measures, such as Nested CVaR (NCVaR), applies the risk measure in each level of the dynamic program independently and preserves most of the favorable computational properties of risk-neutral MDPs Ahmadi et al. (2021a). Unfortunately, nested risk measures are difficult to interpret; their value depends on the sequence in which the rewards are obtained in a complex and unpredictable way Kupper and Schachermayer (2006) and may be unbounded even if MDPs are transient.
While we are unaware of prior work on the TRC objective with ERM or EVaR risk-aversion allowing both positive and negative rewards, the ERM risk measure is closely related to exponential utility functions. Prior work on TRC with exponential utility functions also imposes constraints on the sign of the instantaneous rewards, such as all positive rewards Blackwell (1967) or all negative rewards Bertsekas and Tsitsiklis (1991); Freire and Delgado (2016); Carpin et al. (2016); de Freitas et al. (2020); Fei et al. (2021a, b); Ahmadi et al. (2021a); Cohen et al. (2021); Meggendorfer (2022). Disallowing a mix of positive and negative rewards limits the modeling power of prior work because it requires that either all states are more desirable or all states are less desirable than the terminal state. Allowing rewards with mixed signs raises some technical challenges, which we address by employing a squeeze argument that takes advantage of MDP’s transience.
Notation. We use a tilde to mark random variables, e.g. $\tilde{x}$. Bold lower-case letters represent vectors, and upper-case bold letters represent matrices. Sets are either calligraphic or upper-case Greek letters. The symbol $\mathbb{X}$ represents the space of real-valued random variables. When a function is defined over an index set, such as $z\colon\{1,2,\dots,N\}\to\mathbb{R}$, we also treat it interchangeably as a vector $\bm{z}\in\mathbb{R}^{n}$ such that $z_{i}=z(i),\forall i=1,\dots,n$. Finally, $\mathbb{R},\mathbb{R}_{+},\mathbb{R}_{++}$ denote real, non-negative real, and positive real numbers, respectively. $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. Given a finite set $\mathcal{Y}$, the probability simplex is $\Delta_{\mathcal{Y}}:=\{x\in\mathbb{R}_{+}^{\mathcal{Y}}\mid\bm{1}^{\mathrm{T}}x=1\}$.
@@ 2 Background on risk-averse MDPs
We focus on solving Markov decision processes (MDPs) Puterman (2005), modeled by a tuple $(\bar{\mathcal{S}},\mathcal{A},\bar{p},\bar{r},\bm{\bar{\mu}})$, where $\bar{\mathcal{S}}=\{1,2,\dots,S,S+1\}$ is the finite set of states and $\mathcal{A}=\{1,2,\ldots,A\}$ is the finite set of actions. The transition function $\bar{p}\colon\bar{\mathcal{S}}\times\mathcal{A}\to\Delta_{\bar{\mathcal{S}}}$ represents the probability $\bar{p}(s,a,s^{\prime})$ of transitioning to $s^{\prime}\in\bar{\mathcal{S}}$ after taking $a\in\mathcal{A}$ in $s\in\bar{\mathcal{S}}$ and $\bm{\bar{p}}_{sa}\in\Delta_{\bar{\mathcal{S}}}$ is such that $(\bm{\bar{p}}_{sa})_{s^{\prime}}=\bar{p}(s,a,s^{\prime})$. The function $\bar{r}\colon\bar{\mathcal{S}}\times\mathcal{A}\times\bar{\mathcal{S}}\to\mathbb{R}$ represents the reward $\bar{r}(s,a,s^{\prime})\in\mathbb{R}$ associated with transitioning from $s\in\bar{\mathcal{S}}$ and $a\in\mathcal{A}$ to $s^{\prime}\in\bar{\mathcal{S}}$. The vector $\bm{\bar{\mu}}\in\Delta_{\bar{\mathcal{S}}}$ is the initial state distribution.
@@
The solution to an MDP is a policy. Given a horizon $t\in\mathbb{N}$, a history-dependent policy in the set $\Pi_{\mathrm{HR}}^{t}$ maps the history of states and actions to a distribution over actions. A Markov policy $\pi\in\Pi_{\mathrm{MR}}^{t}$ is a sequence of decision rules $\pi=(\bm{d}_{0},\bm{d}_{1},\dots,\bm{d}_{t-1})$ with $\bm{d}_{k}\colon\mathcal{S}\to\Delta_{\mathcal{A}}$ the decision rule for taking actions at time $k$. The set of all randomized decision rules is $\mathcal{D}=(\Delta_{\mathcal{A}})^{\mathcal{S}}$. Stationary policies $\Pi_{\mathrm{SR}}$ are Markov policies with $\pi:=(\bm{d})_{\infty}:=(\bm{d},\bm{d},\dots)$ with the identical decision rule in every timestep. We treat decision rules and stationary policies interchangeably. The sets of deterministic Markov and stationary policies are denoted by $\Pi_{\mathrm{MD}}^{t}$ and $\Pi_{\mathrm{SD}}$. Finally, we omit the superscript $t$ to indicate infinite horizon definitions of policies.
@@
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\mathbb{E}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right],$ (1)
@@
Unlike the discounted criterion, the risk-neutral TRC may be unbounded, optimal policies may not exist, or may be non-stationary Bertsekas and Yu (2013); James and Collins (2006). To circumvent these issues, we assume that all policies have a positive probability of eventually transitioning to the sink state.
@@
2.2 underlies most of our results. Transient MDPs are important because their optimal policies exist and can be chosen to be stationary deterministic (Kallenberg, 2021, theorem 4.12). Transient MDPs are also common in stochastic games Filar and Vrieze (2012) and generalize the stochastic shortest path problem Bertsekas and Yu (2013).
An important tool in their analysis is the spectral radius $\rho\colon\mathbb{R}^{n\times n}\to\mathbb{R}$ which is defined for each $\bm{A}\in\mathbb{R}^{n\times n}$ as the maximum absolute eigenvalue: $\rho(\bm{A}):=\max_{i=1,\dots,n}|\lambda_{i}|$ where $\lambda_{i}$ is the $i$-th eigenvalue Horn and Johnson (2013).
@@
An MDP is transient if and only if $\rho(\bm{P}^{\pi})<1$ for all $\pi\in\Pi_{\mathrm{SR}}$.
Now, let us understand the differences between a discounted MDP and a transient MDP, which are useful in demonstrating the behavior of risk-averse objectives. Consider the MDPs in Figure 1. There is one non-sink state $s$ and one action $a$. A triple tuple represents an action, transition probability, and a reward separately. Note that every discounted MDP can be converted to a transient MDP as described in Su et al. (2024a, appendix B). For the discounted MDP, the discount factor is $\gamma$. For the transient MDP, $e$ is the sink state, and there is a positive probability $1-\epsilon$ of transiting from state $s$ to state $e$. Once the agent reaches the state $e$, it stays in $e$. For the risk-neutral objective, if $\gamma$ equals $\epsilon$, their value functions have identical values. However, for risk-aversion objectives, such as ERM, we show that the value functions in a discounted MDP can diverge from those in a transient MDP in Section 5.
@@
Monetary risk measures aim to generalize the expectation operator to account for the spread of the random variable. Entropic risk measure (ERM) is a popular risk measure, defined for any risk level $\beta>0$ and $\tilde{x}\in\mathbb{X}$ as Follmer and Schied (2016)
@@
and extended to $\beta\in[0,\infty]$ as $\operatorname{ERM}_{0}[\tilde{x}]=\lim_{\beta\to 0^{+}}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\mathbb{E}[\tilde{x}]$ and $\operatorname{ERM}_{\infty}[\tilde{x}]=\lim_{\beta\to\infty}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$. ERM plays a unique role in sequential decision-making because it is the only law-invariant risk measure that satisfies the tower property (e.g., Su et al. (2024a, proposition A.1)), which is essential in constructing dynamic programs Hau et al. (2023b). Unfortunately, two significant limitations of ERM hinder its practical applications. First, ERM is not positively homogenous and, therefore, the risk value depends on the scale of the rewards, and ERM is not coherent Follmer and Schied (2016); Hau et al. (2023b); Ahmadi-Javid (2012). Second, the risk parameter $\beta$ is challenging to interpret and does not relate well to other standard risk measures, like VaR or CVaR.
@@
and extended to $\operatorname{EVaR}_{0}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$ and $\operatorname{EVaR}_{1}\left[\tilde{x}\right]=\mathbb{E}\left[\tilde{x}\right]$ Ahmadi-Javid (2012). It is important to note that the supremum in (4) may not be attained even when $\tilde{x}$ is a finite discrete random variable Ahmadi-Javid and Pichler (2017).
EVaR addresses the limitations of ERM while preserving its benefits. EVaR is coherent and positively homogenous. EVaR is also a good approximation to interpretable quantile-based risk measures, like VaR and CVaR Ahmadi-Javid (2012); Hau et al. (2023b).
@@
Risk-averse MDPs, using static VaR and CVaR risk measures, under the discounted criterion received abundant attention Hau et al. (2023a); Bäuerle and Ott (2011); Bäuerle and Glauner (2022); Pflug and Pichler (2016); Li et al. (2022), showing that these objectives require history-dependent optimal policies. In contrast, nested risk measures under the TRC may admit stationary policies that can be computed using dynamic programming Ahmadi et al. (2021a); Meggendorfer (2022); de Freitas et al. (2020); Gavriel et al. (2012). However, the TRC with nested CVaR can be unbounded (Su et al., 2024a, proposition C.1). Recent work has shown that optimal Markov policies exist for EVaR discounted objectives, and they can be computed via dynamic programming Hau et al. (2023b), building upon similar results established for ERM Chung and Sobel (1987). However, in TRC with ERM, the value functions may also be unbounded (Su et al., 2024a, proposition D.1).